OpenStudy (anonymous):

A town has a population of 100,000. The population increases in a geometric sequence by 3% (0.03) each year. What is the population in 15 years?

OpenStudy (anonymous):

Formula r=t1/tn-1 tn=t1r^n-1

OpenStudy (anonymous):

1 year later, population = 100,000 + 0.03*100,000 = 100,000*(1 + 0.03) 2 years later, population = (100,000*(1+ 0.03)) + 0,03*(100,000*(1 + 0.03)) = 100,000*(1 + 0,03)*(1 + 0,03) = 100,000*(1+0,03)^2 and so on so, the formula of the population for n years later = 100,000*(1 + 0.03)^n 15 years later, substitute n = 15, and you'll get the answer

OpenStudy (anonymous):

So 100,000(1 + 0.03)^15

OpenStudy (anonymous):

155797. @alicealc

OpenStudy (anonymous):

yes